OpenStudy (loser66):

Statistics help, please. An insurance company estimates that 40% of policyholders who have only an auto policy will renew next year and 60% of policyholders who have only a homeowners policy will renew next year. The company estimates that 80% of policyholders who have both an auto and a homeowners policy will renew at least one of those policies next year. Company records show that 65% of policyholders have an auto policy, 50% of policyholders have a homeowners policy, and 15% of policyholders have both an auto and a homeowners policy. Using the company's estimates ..(continue on comments)

OpenStudy (loser66):

calculate the percentage of policyholders that will renew at least one policy next year. Options: 20 29 53 70

OpenStudy (mathmate):

This is a trick question! When you add up the percentage of clients, you get 65+50+15=130% which is impossible. However, when we realize that the 15% is included in both the 65% (auto) and 50% (home) policies, then things will make sense. So we have auto=65-15=50% home=50-15=35% both=15% Total = 50+35+15=100% Can you finish the rest, using law of total probabilities?

OpenStudy (loser66):

I got it. Thanks @mathmate